A double-slit interference pattern is created by two narrow slits spaced 0.22 mm apart. The distance between the first and the fifth minimum on a screen 58 cm behind the slits is 6.3 mm.

What is the wavelength (in nm) of the light used in this experiment?
Answer #1

distance between 1st 4 sth mimm - 98xL - AxL = 6.3 mm 2d zd 4x2x0.58 = 6.3x103 0.22x10-3 - > 2= 597-4 nm Answer =